OpenStudy (anonymous):

find the roots -x3 + 5x2 - 11x + 55 = 0

OpenStudy (anonymous):

@jim_thompson5910 can you help me?

jimthompson5910 (jim_thompson5910):

hint: factor by grouping

OpenStudy (anonymous):

im not sure how to do that

jimthompson5910 (jim_thompson5910):

-x^3 + 5x^2 - 11x + 55 = 0 (-x^3 + 5x^2) + (-11x + 55) = 0 -x^2(x - 5) -11(x - 5) = 0 I'll let you finish up

OpenStudy (abb0t):

you could also use the fundamental theorem of algebra, using long division

OpenStudy (anonymous):

okay I factored it then what?

jimthompson5910 (jim_thompson5910):

what did you get when you factored?

OpenStudy (anonymous):